# The inclusive Synthetic Control Method

### v1

@@ 8 Conclusion
We introduce iSCM, a modification of the SCM that allows the inclusion of units “potentially affected” by an intervention in the donor pool. Our method is useful in applications where it is either important to include other treated units in the donor pool or where some of the units in the donor pool are affected indirectly by the intervention (spillover effects). Our iSCM requires that we choose a priori which units are “potentially affected” by the treatment and that the assumptions of the chosen SC-type estimator are valid. In addition, our methods can only be used if at least one “pure” control unit receives non-zero weight. A major advantage of iSCM is that it can be easily implemented using the synthetic control estimator or many of the new estimation methods available in the literature. Moreover, we demonstrate that iSCM is almost certain to enhance the performance compared to the “unrestricted” SCM. Additionally, we introduce a data-driven approach to determine whether iSCM could potentially outperform the “restricted” SCM. Even in situations where excluding “potentially affected” units from the donor pool does not seem to be harmful, our iSCM can serve as a robustness check. Finally, we illustrate the use of iSCM by re-estimating the impact of German reunification on GDP per capita, confirming Abadie et al. (2015) expectations about the potential direction of the spillover effect from West Germany to Austria. We find small negative spillover effects on Austria, which would imply an even more negative treatment effect on West Germany.
@@ A.1 Non-singularity
A.2 A geometrical interpretation on why use iSCM
In Figures 3 and 4, we graphically represent possible scenarios from the point of view of unit 1 (“main treated” unit) on the left side and unit 2 (“potentially affected” unit) on the right side. Without loss of generality, we assume to observe only two predictors ($x_{1}$ and $x_{2}$) for each unit. $X_{1}$, i.e., the red point, is the vector that includes the pre-intervention predictors of the “main treated” and $X_{2}$, i.e., the blue point, is the vector that includes the pre-intervention predictors of the only affected unit. All other points represent the vectors of pre-intervention predictors of each “pure control” unit. When marked in black, they contribute to the synthetic control, whereas when marked in grey they do not contribute.
Figure 3 shows the scenario in which both the “main treated” and the “potentially affected” units lie inside the convex hull of the “pure control” units. In this case, one can reproduce $X_{1}$ only using “pure control” units. However, a closer look at the right side of Panel 3a reveals that including the “potentially affected” unit to reproduce the characteristics of the “main treated” unit allows the exclusion of the farthest “pure control” unit (unit 3), restricting the donor pool, and potentially reducing interpolation bias (see Abadie 2021). The same goes for the “potentially affected” unit: including the “main treated” in the donor pool allows the exclusion of unit 4, which lies farther away from unit 2.
@@ A.2 A geometrical interpretation on why use iSCM
Figure 4 presents some scenarios in which only one unit between the “main treated” and the “potentially affected” unit lies inside the convex hull of the other and the “pure control” units. The left side of Panel 4a shows the case in which the “main treated” unit lies outside but close to the convex hull. Regardless of whether we exclude or include the “potentially affected” unit, we can only approximate $X_{1}$. However, excluding unit 2 would lead to a bigger discrepancy between $X_{1}-X_{0}W$, therefore it is better to include it in the donor pool. On the right side of Panel 4a, we notice that the “potentially affected” unit lies outside the convex hull unless we include the “main treated” in its donor pool. Panel 4b shows a symmetric situation to that in Panel 4a. The left side of Panel 4c shows a scenario in which the “main treated” lies outside the convex hull and including unit 2 improves the approximation. The right side of Panel 4c shows a scenario in which unit 2 is always in the convex hull and including unit 1 the convex hull becomes bigger but it allows to improve approximation because unit 3 can be excluded. Panel 4d describes a scenario in which using iSCM clearly makes things worse. We suggest using iSCM when the “potentially affected” units receive substantial weight. It is less likely in this scenario.

### v2

@@ 8 Conclusion
We introduce iSCM, a modification of the SCM that allows the inclusion of units ”potentially affected” by an intervention in the donor pool. Our method is useful in applications where it is either important to include other treated units in the donor pool or where some of the units in the donor pool are affected indirectly by the intervention (spillover effects). Our iSCM requires that we choose a priori which units are ”potentially affected” by the treatment and that the assumptions of the chosen SC-type estimator are valid. In addition, our methods can only be used if at least one “pure” control unit receives non-zero weight. A major advantage of iSCM is that it can be easily implemented using the synthetic control estimator or many of the new estimation methods available in the literature. Moreover, we demonstrate that iSCM is almost certain to enhance the performance compared to the ”unrestricted” SCM. Additionally, we introduce a data-driven approach to determine whether iSCM could potentially outperform the ”restricted” SCM. Even in situations where excluding ”potentially affected” units from the donor pool does not seem to be harmful, our iSCM can serve as a robustness check. Finally, we illustrate the use of iSCM by re-estimating the impact of German reunification on GDP per capita. Using both the original and penalized SCM, we confirm Abadie et al. (2015) expectations about the direction of the spillover effect from West Germany to Austria, findings small negative spillover effects on Austria. This implies that the negative treatment effect on West Germany might be larger than previously estimated.
@@ A.1 Non-singularity
A.2 RMSPEs ratios
Panels 3a and 3b of Figure 3 show the ratios between the RMSPEs in the post- and pre-reunification of West Germany’s donor pool and Austria’s donor pool, respectively. We can observe that West Germany’s value is very high and is the largest compared to any other countries in the donor pool. On the other hand, Austria’s RMSPE ratio is the second lowest indicating the spillover effect is not significant.
(a) RMSPE ratios West Germany
(b) RMSPE ratios Austria
A.3 A geometrical interpretation on why use iSCM
In Figures 4 and 5, we graphically represent possible scenarios from the point of view of unit 1 (”main treated” unit) on the left side and unit 2 (”potentially affected” unit) on the right side. Without loss of generality, we assume to observe only two predictors ($x_{1}$ and $x_{2}$) for each unit. $X_{1}$, i.e., the red point, is the vector that includes the pre-intervention predictors of the ”main treated” and $X_{2}$, i.e., the blue point, is the vector that includes the pre-intervention predictors of the only affected unit. All other points represent the vectors of pre-intervention predictors of each ”pure control” unit. When marked in black, they contribute to the synthetic control, whereas when marked in grey they do not contribute.
Figure 4 shows the scenario in which both the ”main treated” and the ”potentially affected” units lie inside the convex hull of the ”pure control” units. In this case, one can reproduce $X_{1}$ only using ”pure control” units. However, a closer look at the right side of Panel 4a reveals that including the ”potentially affected” unit to reproduce the characteristics of the ”main treated” unit allows the exclusion of the farthest ”pure control” unit (unit 3), restricting the donor pool, and potentially reducing interpolation bias (see Abadie 2021). The same goes for the ”potentially affected” unit: including the ”main treated” in the donor pool allows the exclusion of unit 4, which lies farther away from unit 2.
@@ A.3 A geometrical interpretation on why use iSCM
Figure 5 presents some scenarios in which only one unit between the ”main treated” and the ”potentially affected” unit lies inside the convex hull of the other and the ”pure control” units. The left side of Panel 5a shows the case in which the ”main treated” unit lies outside but close to the convex hull. Regardless of whether we exclude or include the ”potentially affected” unit, we can only approximate $X_{1}$. However, excluding unit 2 would lead to a bigger discrepancy between $X_{1}-X_{0}W$, therefore it is better to include it in the donor pool. On the right side of Panel 5a, we notice that the ”potentially affected” unit lies outside the convex hull unless we include the ”main treated” in its donor pool. Panel 5b shows a symmetric situation to that in Panel 5a. The left side of Panel 5c shows a scenario in which the ”main treated” lies outside the convex hull and including unit 2 improves the approximation. The right side of Panel 5c shows a scenario in which unit 2 is always in the convex hull and including unit 1 the convex hull becomes bigger but it allows to improve approximation because unit 3 can be excluded. Panel 5d describes a scenario in which using iSCM clearly makes things worse. We suggest using iSCM when the ”potentially affected” units receive substantial weight. It is less likely in this scenario.